OpenStudy (anonymous):

Which of the following ordered pairs is a solution to the inequality y less than or greater to -2x + 8? (0, -4) (5, -1) (7, 0) (9, 1)

OpenStudy (anonymous):

y <= -2x + 8

OpenStudy (texaschic101):

y <= -2x + 8 lets start subbing.... (0,-4) -4 <= -2(0) + 8 -4 <= 0 + 8 -4 <= 8 (true) (5,-1) -1 <= -2(5) + 8 -1 <= -10 + 8 -1 <= -2 (false) (7,0) 0 <= -2(7) + 8 0 <= -14 + 8 0 <= - 6 (false) (9,1) 1 <= -2(9) + 8 1 <= -18 + 8 1 <= - 10 (false)

OpenStudy (texaschic101):

do you understand how I did that ?

OpenStudy (texaschic101):

In a solution set, the first number is x and the second number is y....(x,y) I just subbed the number into the inequality and found which one was true.

OpenStudy (anonymous):

Yeah sorry i didnt see your comment

OpenStudy (anonymous):

did you find the awnser??